# Observer-robust energy condition verification for warp drive spacetimes

### 03 Mar 2026 v2

@@ Abstract
We present warpax, an open-source, GPU-accelerated Python toolkit for observer-robust energy condition analysis of warp drive spacetimes. Existing tools evaluate energy conditions for a finite sample of observer directions; warpax replaces discrete sampling with continuous, gradient-based optimization over the timelike observer manifold (rapidity and boost direction), backed by Hawking–Ellis algebraic classification. At Type I stress-energy points, which comprise > 96{>}\,96% of all grid points across the tested metrics, an algebraic eigenvalue check determines energy-condition satisfaction exactly, independent of any observer search or rapidity cap. At non-Type I points the optimizer provides rapidity-capped diagnostics. Stress-energy tensors are computed from the ADM metric via forward-mode automatic differentiation, eliminating finite-difference truncation error. Geodesic integration with tidal-force and blueshift analysis is also included.
We analyze five warp drive metrics (Alcubierre, Lentz, Van Den Broeck, Natário, Rodal) and one warp shell metric (used primarily as a numerical stress test). For the Rodal metric, the standard Eulerian-frame analysis misses violations at over $28\%$ of grid points (dominant energy condition) and over $15\%$ (weak energy condition). Even where the Eulerian frame identifies the correct violation set, observer optimization reveals that violation severity can be orders of magnitude larger (e.g. Alcubierre weak energy condition:  ${\sim}\,90{,}000\times$ at rapidity cap $\zeta_{\max}=5$, scaling as $e^{2\zeta_{\max}}$; the rapidity cap is an analysis hyperparameter, not a physically preferred bound). These results demonstrate that single-frame evaluation can systematically underestimate both the spatial extent and the magnitude of energy condition violations in warp drive spacetimes. warpax is freely available at https://github.com/anindex/warpax.
@@ 1 Introduction
The WarpFactory toolkit [4] addresses this by evaluating energy conditions over a discrete set of sampled observer velocities and directions (typically $\sim$1000 null and timelike vectors), taking the minimum across all samples at each grid point. This sampling-based approach is a significant advance over single-frame evaluation, but has inherent limitations: the accuracy depends on the sampling density, and the discrete set may miss narrow violation cones in the observer manifold. WarpFactory also relies on fourth-order central finite differences for the curvature computation [5], which introduces truncation error and requires a step-size parameter.
The observer-dependence of energy condition evaluation is well-established in the mathematical relativity literature. Hawking and Ellis [3] classified the algebraic types of the stress-energy tensor (Types I–IV), and the conditions under which each energy condition can be violated depend on this algebraic structure. For Type I (the generic case), the NEC is satisfied for all observers if and only if $\rho+p_{i}\geq 0$ for each principal pressure $p_{i}$, but evaluating this requires computing the eigenvalues of the mixed tensor $T^{a}{}_{b}$, rather than simply reading off components in a coordinate basis.
In this paper, we present warpax, a JAX-based toolkit that replaces discrete observer sampling with continuous, gradient-based observer optimization over the observer manifold. warpax is built on JAX [6] for GPU-accelerated automatic differentiation, uses Equinox [7] for module structure, and provides:
@@
We analyze five warp drive metrics and one warp shell geometry. For the Rodal metric (we use a lab-frame standardization $F=1-f_{\rm paper}$, $G=1-g_{\rm paper}$; see Appendix D), the Eulerian single-frame DEC analysis fails to detect violations at over 28% of all grid points; for WarpShell (a regularized implementation used primarily as a numerical stress test; see Section 7.3), missed WEC violations are below $0.1\%$ of the grid. For Alcubierre and Natário, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis (missed =0=0 at the sampled resolution); Lentz shows zero missed NEC/WEC points but nonzero SEC misses, and is severely under-resolved ($\ll\!1$ cell across the wall at $50^{3}$); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$ at $v_{s}=0.5$). The $\zeta_{\max}$-capped violation severity (at $\zeta_{\max}=5$) can be substantially larger in regions already identified as violating. We validate against WarpFactory’s published Eulerian results and quantify the additional violations discovered by observer optimization where they exist.
Related work.
@@
Computational tools for general relativity span a broad spectrum from symbolic computer algebra to large-scale numerical evolution (see Ref. [8] for a comprehensive survey). Symbolic packages such as xAct [9] for Mathematica and SageManifolds [10] for SageMath provide exact tensor calculus on differentiable manifolds, enabling analytic derivation of curvature tensors and energy condition expressions in closed form. These tools are invaluable for deriving exact results but do not perform numerical verification of energy conditions on discretized spacetime grids or systematic optimization over observer families. At the numerical end, community frameworks for numerical relativity (e.g. the Einstein Toolkit, NRPy+) focus on dynamical evolution of Einstein’s equations rather than pointwise energy condition assessment of static metrics. The use of automatic differentiation for computing curvature tensors in general relativity is an emerging paradigm: libraries such as diffjeom [11] demonstrate JAX-based computation of the Christoffel–Riemann–Ricci hierarchy, and Bara [12] identifies solving field equations in exotic spacetimes as an underexplored domain for gradient-based methods. To the best of our knowledge, warpax is the first tool to combine exact autodiff-derived curvature with systematic, gradient-based optimization over the timelike observer manifold for energy condition verification.
@@
The paper is organized as follows. Section 2 presents the mathematical methods. Section 3 describes the warpax toolkit. Section 4 presents energy condition results. Section 5 presents geodesic results. Section 6 analyzes kinematic scalars. Section 7 validates numerical convergence. Section 8 discusses implications, and Section 9 concludes. Appendices provide supplementary material: transition-function ablation (Appendix A), Rodal DEC ablation (Appendix B), wall-resolution evidence (Appendix C), and full metric definitions (Appendix D).
@@ 2.2 Hawking–Ellis classification
  • Type II: A double null eigenvector (radiation-like). Stress-energy has the form of a null fluid plus a Type I part.
  • Type III: A triple null eigenvector (non-generic nilpotent stress–energy). Rare in practice.
@@
Throughout this paper, “robust margin” refers to the merged quantity (hereafter hybrid margin): algebraic slack at Type I points and $\zeta_{\max}$-capped extremum at non-Type I points. For Type I points, violation detection and the reported margins use the algebraic slack (item 2), which is cap-independent and authoritative. The observer optimizer is run additionally to provide the explicit worst-case observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$; its $\zeta_{\max}$-capped contraction minimum (item 3) is reported separately in sensitivity analyses (Table 9) but is not used for violation detection. The algebraic slack and capped extremum are in general different numerical values: the former represents the global eigenvalue-based quantity, while the latter depends on $\zeta_{\max}$. These two quantities are conceptually distinct and should not be compared directly; the algebraic slack measures distance to the violation boundary, while the capped extremum quantifies the worst-case energy density within the tested observer range. The algebraic slack (e.g. $\min_{i}(\rho+p_{i})$ for NEC) certifies satisfaction or violation but is not, in general, equal to the minimum observer-contracted energy density $\min_{u}T_{ab}\,u^{a}u^{b}$; it measures proximity to the eigenvalue inequality boundary rather than a physical extremum over the observer manifold. For the metrics and rapidity caps used here ($\zeta_{\max}=5$) we empirically observe sign agreement between the algebraic Type I truth and the $\zeta_{\max}$-capped optimizer at all grid points (i.e. no violation is “hidden” above the cap). This is not guaranteed in general: a Type I tensor with $\rho+p_{i}<0$ but $|\rho+p_{i}|\ll|\rho|$ could require an arbitrarily large boost to flip the sign of the measured energy density, so a finite $\zeta_{\max}$ might in principle miss it. At non-Type I grid points (up to $\sim\!3$% for the warp metrics and <1<\!1% for WarpShell; see Table 8), the algebraic inequalities do not apply and the optimizer is the sole method; this pathway is validated via a null-dust benchmark (Section 8).
@@ 2.3 Observer parameterization
To verify energy conditions over all admissible observers, we parameterize the space of timelike and null vectors at each point.
@@
The rapidity parameterization has the advantage that $\zeta=0$ recovers the Eulerian observer, so the optimization landscape always includes the Eulerian result as a baseline. The WEC and DEC involve three-dimensional optimization over $(\zeta,\theta,\phi)$, while the NEC requires only $(\theta,\phi)$. The true timelike observer manifold has unbounded rapidity ($\zeta\in[0,\infty)$); in practice, we optimize over $\zeta\in[0,\zeta_{\max}]$ with a fixed cap $\zeta_{\max}=5$ (corresponding to $\gamma\approx 74$; see Section 2.4).

### 27 Apr 2026 v3

@@ Abstract
We present warpax, an open-source, GPU-accelerated Python toolkit for observer-robust energy-condition verification of warp drive spacetimes, together with a benchmark application to six warp-drive geometries that demonstrates the methodology and produces new quantitative findings.
Existing tools evaluate energy conditions for a finite sample of observer directions. warpax replaces discrete sampling with continuous, gradient-based optimization over the full timelike observer manifold, backed by Hawking–Ellis algebraic classification. At Type I stress-energy points—which dominate all tested metrics—an algebraic eigenvalue check determines energy-condition satisfaction exactly, independent of any observer search. At non-Type I points the optimizer provides rapidity-capped diagnostics. Stress-energy tensors are computed from the Arnowitt–Deser–Misner metric via forward-mode automatic differentiation, eliminating finite-difference truncation error. Geodesic integration with tidal-force and blueshift analysis is also included.
We apply warpax to five warp drive metrics (Alcubierre, Lentz, Van Den Broeck, Natário, Rodal) and one warp shell metric. For several metrics, the standard Eulerian-frame analysis misses a significant fraction of violated grid points entirely; even where it identifies the correct violation set, observer optimization reveals that violation magnitudes can be orders of magnitude larger. These results demonstrate that single-frame evaluation can systematically underestimate both the spatial extent and the severity of energy-condition violations in warp drive spacetimes. Throughout, we distinguish the invariant energy density (eigenvalue of $T^{a}{}_{b}$) from the observer-dependent $T_{ab}\,u^{a}u^{b}$ and the Eulerian projection that appears in prior work.
All results use subluminal bubble velocities; at superluminal speeds the Alcubierre-family metrics develop signature changes that fall outside our current assumptions. warpax is freely available at https://github.com/anindex/warpax.
@@ 1 Introduction
We make two co-equal contributions. The first is a toolkit: we present warpax, an open-source JAX-based Python package that replaces discrete observer sampling with continuous, gradient-based observer optimization over the observer manifold. warpax is built on JAX [7] for GPU-accelerated automatic differentiation, uses Equinox [22] for module structure, and provides:
@@
Our second contribution is a benchmark application of the toolkit to six warp-drive metrics that demonstrates the methodology and produces the quantitative findings reported below.
WarpFactory [21] addresses observer coverage by evaluating energy conditions over a discrete set of sampled observer velocities and directions (typically $\sim$1000 null and timelike vectors), taking the minimum across all samples at each grid point. This sampling-based approach is a significant advance over single-frame evaluation, but has inherent limitations: the accuracy depends on the sampling density, and the discrete set may miss narrow violation cones in the observer manifold. WarpFactory also relies on fourth-order central finite differences for the curvature computation [20], which introduces truncation error and requires a step-size parameter.
@@
The observer-dependence of energy condition evaluation is well-established in the mathematical relativity literature. Hawking and Ellis [19] classified the algebraic types of the stress-energy tensor (Types I–IV), and the conditions under which each energy condition can be violated depend on this algebraic structure. For Type I (the generic case), the NEC is satisfied for all observers if and only if $\rho+p_{i}\geq 0$ for each principal pressure $p_{i}$, but evaluating this requires computing the eigenvalues of the mixed tensor $T^{a}{}_{b}$, rather than simply reading off components in a coordinate basis. We analyze warp drive metrics in the subluminal regime $v_{s}<1$. At $v_{s}=1$ the Alcubierre- and Lentz-type metrics develop vanishing $g_{00}$ along the bubble wall (equivalently $\det g\to 0$), and for $v_{s}>1$ the signature flips inside the wall region ($g_{00}>0$) so the affected region is no longer a Lorentzian spacetime. Our observer-optimization framework assumes a globally Lorentzian background; accordingly, all results reported here use $v_{s}\leq 0.99$. See Section 1.2 for a fuller discussion of scope and limitations. We analyze five warp drive metrics and one warp shell geometry. For the Rodal metric (we use a lab-frame standardization $F=1-f_{\rm paper}$, $G=1-g_{\rm paper}$; see Appendix D), the Eulerian single-frame DEC analysis fails to detect violations at over 28% of all grid points; for WarpShell (a regularized implementation used primarily as a numerical stress test; see Section 7.3), missed WEC violations are below $0.1\%$ of the grid. For Alcubierre and Natário, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis (missed =0=0 at the sampled resolution); Lentz shows zero missed NEC/WEC points but nonzero SEC misses, and is severely under-resolved ($\ll\!1$ cell across the wall at $50^{3}$); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$ at $v_{s}=0.5$). The $\zeta_{\max}$-capped violation severity (at $\zeta_{\max}=5$) can be substantially larger in regions already identified as violating. We validate against WarpFactory’s published Eulerian results and quantify the additional violations discovered by observer optimization where they exist.
1.1 Related work
The literature relevant to this paper falls into three strands: (i) energy-condition analysis in warp-drive spacetimes, (ii) the 2021–2026 proliferation of warp-drive geometries, and (iii) numerical methods for general relativity and energy-condition verification. We survey each in turn.
1.1.1 Energy-condition analysis in warp-drive spacetimes.
The observer-dependence of energy-condition evaluation is well established in the mathematical-relativity literature. Hawking and Ellis [19] introduced the algebraic classification of $T^{a}{}_{b}$ into four types; Martín-Moruno and Visser have since provided a modern treatment of the classification and of the conditions under which each energy condition can be violated for each type, both in the “essential core” formulation [30] and in a dedicated analysis of the rare Type III case [29]. The generalized Rainich conditions [28] and the test-field versus back-reaction distinction [31] extend the Hawking–Ellis framework to the stress-energy conditions that motivate our $T_{ab}\,u^{a}u^{b}$ optimization.
Santiago, Schuster, and Visser [40] proved that any physically reasonable warp drive must violate the null energy condition (NEC), a universal statement that does not depend on the choice of observer. Celmaster and Rubin [8] subsequently demonstrated explicit Eulerian-frame weak-energy-condition (WEC) violations for the Lentz geometry, showing that the original positive-energy claim is incorrect even in the single preferred frame. In parallel, the Bobrick–Martire taxonomy of physical warp drives [6] and the hidden-geometric positive-energy construction of Fell and Heisenberg [13] reframe the problem from “violations in a single metric” to “which warp-drive class admits a physically acceptable stress-energy”. Our contribution to this strand is an observer-robust verification methodology that reports wall-restricted Hawking–Ellis type breakdowns together with NEC/WEC/SEC/DEC margins from continuous observer optimization, augmenting the Eulerian-only diagnostics used in prior work and distinguishing the invariant energy density $\rho$ from the observer-contracted $T_{ab}\,u^{a}u^{b}$ that appears in single-frame analysis.
1.1.2 Warp-drive geometries.
The foundational warp-drive constructions remain Alcubierre’s 1994 metric [1], Natário’s zero-expansion variant [32], and Van Den Broeck’s modified-volume construction [43]. The Lentz soliton [25] was proposed as a positive-energy alternative but was subsequently shown to violate the NEC [40] and the WEC in the Eulerian frame [8]. The Rodal metric [38] completes our benchmark set; Rodal’s own earlier work [35, 36] analyses curvature invariants for Alcubierre and Natário using the same 3+1 formalism that underpins our ADM setup.
Recent 2024–2026 constructions enlarge this space considerably. Fuchs et al. [15] introduce a constant-velocity shell-based positive-energy warp drive in the Bobrick–Martire family. Garattini and Zatrimaylov [16] show that embedding a warp drive on a black-hole background alleviates energy-condition violations, and further generalize to a de Sitter background [17]. Clough, Dietrich, and Khan [9] report the first numerical evolution of a warp-drive collapse and the associated gravitational-wave signature, the dynamical counterpart to the static analysis we perform here. Rodal [37, 39] provides theoretical arguments against low-energy and fast Type I warp drives via metamaterial-coupling and weakly-birefringent-screening considerations. Santos-Pereira et al. [41] study Darmois junction conditions for matching Alcubierre to Minkowski, a theme related to the shell-regularization used in our WarpShell stress-test geometry. Barzegar, Buchert, and Vigneron [5, 4] provide a modern formal classification of the current warp-drive spacetimes, identifying shared restrictions (flow-orthogonal foliations, vanishing spatial curvature) and possible routes to improvement. These recent contributions motivate the multi-metric, methodology-forward approach to energy-condition analysis that we adopt here.
1.1.3 Numerical methods for general relativity and energy-condition verification.
@@ 1.1 Related work
Computational tools for general relativity span a broad spectrum from symbolic computer algebra to large-scale numerical evolution (see Ref. [26] for a broad survey). Symbolic packages such as xAct [27] for Mathematica and SageManifolds [18] for SageMath provide exact tensor calculus on differentiable manifolds, enabling analytic derivation of curvature tensors and energy-condition expressions in closed form. These tools are invaluable for deriving exact results but do not perform numerical verification of energy conditions on discretized spacetime grids or systematic optimization over observer families. At the numerical end, community frameworks for numerical relativity, in particular the Einstein Toolkit [12], focus on dynamical evolution of Einstein’s equations rather than pointwise energy-condition assessment of static metrics. Neural-network compression of 4D numerical simulations [11] offers an orthogonal axis of computational compression that complements, rather than replaces, pointwise methodology.
The warp-drive community has a dedicated toolkit, WarpFactory, described in a methodology companion [20], analyzed in detail for warp-drive spacetimes by Helmerich et al. [21], and documented online [2]. WarpFactory evaluates energy conditions over a discrete sample of $\sim\!10^{3}$ observer directions and velocities and uses fourth-order central finite differences for the curvature chain. Our warpax toolkit differs methodologically on three axes: it uses forward-mode automatic differentiation (implemented in JAX [7]) for the curvature chain, rather than finite differences, following the emerging autodiff-for-GR approach demonstrated by diffjeom [10] and by the broader astrophysical autodiff program that Bara [3] identifies; it replaces discrete observer sampling with continuous, gradient-based optimization using Optimistix [24]; and it integrates geodesics via Diffrax [23] with an Equinox [22] module structure. To the best of our knowledge, warpax is the first tool to combine exact autodiff-derived curvature with systematic, gradient-based optimization over the timelike observer manifold for energy-condition verification.
@@
The rest of this paper is organized as follows. Section 2 presents the mathematical methods. Section 3 describes the warpax toolkit. Section 4 presents energy condition results. Section 5 presents geodesic results. Section 6 analyzes kinematic scalars. Section 7 validates numerical convergence. Section 8 discusses implications, and Section 9 concludes. Appendices provide supplementary material: transition-function ablation (Appendix A), Rodal DEC ablation (Appendix B), wall-resolution evidence (Appendix C), and full metric definitions (Appendix D).
1.2 Scope and limitations
We analyze warp drive metrics in the subluminal regime $v_{s}<1$. At $v_{s}=1$ the Alcubierre and Lentz metrics develop a vanishing $g_{00}$ along the bubble wall (the lapse $\alpha\to 0$, equivalently $\det g\to 0$), and for $v_{s}>1$ the signature flips inside the wall region ($g_{00}>0$) so that the affected region is no longer a Lorentzian spacetime. Our observer-optimization framework assumes a globally Lorentzian background; accordingly, all results reported here use $v_{s}\leq 0.99$. The superluminal extension is both a physical regime change (classical spacetime breakdown at and beyond $v_{s}=1$) and a computational barrier (NaN propagation through autodiff once $g_{00}$ changes sign).
Our energy condition verification is pointwise and classical: we evaluate $T_{ab}$ at grid points and test eigenvalue inequalities or observer-optimized contractions. We do not compute averaged (line-integrated) or semiclassical bounds such as quantum inequalities [14, 34], which are the natural next step for relating pointwise EC violations to physical viability.
Beyond the subluminal-regime scope of the above results, warpax further provides a differentiable shape-function parametrization for the metric’s bubble profile with three basis families (cubic B-spline, Bernstein polynomial, and Gaussian mixture), exposed through a constrained projected-gradient BFGS optimizer (hard-bound projected-gradient BFGS) on the observer-robust EC margin. The parametrization enables end-to-end gradient descent over shape parameters while honoring lapse-floor, global-hyperbolicity, and bubble-finiteness constraints at every step. Numerical results for the shape-function design catalog (Table 15), including recovery of the Alcubierre profile and new single-condition shape families, are presented in an accompanying arXiv note.
GPU-speedup scope caveat.
The curvature-chain evaluation benefits from GPU parallelism at large grid sizes (Section 2 reports order-of-seconds timings on an NVIDIA A100 at $50^{3}$ and $100^{3}$). At the observer-optimization inner-loop scale, however, GPU performance is workload-dependent: a single-point benchmark on an NVIDIA RTX 5090 (Blackwell architecture) reports a geometric-mean BFGS speedup of $0.60\times$ across four optimizer variants, driven by just-in-time (JIT) compile and GPU-transfer overhead at the single-point scale. Full $50^{3}$ / $100^{3}$ end-to-end speedups at field-scale workload are expected to exceed CPU performance significantly but are not directly measured here.
@@ 2.1 Stress-energy via automatic differentiation
Three energy-density quantities.
We distinguish three energy-density quantities that appear throughout this paper. The invariant energy density $\rho$ is defined as the timelike eigenvalue of the mixed stress-energy tensor $T^{a}{}_{b}$ and is frame-independent; it equals $T_{ab}\,u^{a}u^{b}$ when evaluated on the principal timelike eigenvector. The observer-dependent energy density $T_{ab}\,u^{a}u^{b}$ is the energy density measured by an observer with four-velocity $u^{a}$ and depends on the observer choice. The Eulerian energy density $T_{ab}\,n^{a}n^{b}$ is the special case where $n^{a}$ is the unit normal to the ADM spatial hypersurface; it is the single-frame baseline used by prior work that reports “energy density” without qualification.
Notation.
We use $R_{b}$ throughout to denote the warp bubble radius, distinct from the Ricci scalar $R$ that appears in curvature expressions. The shell radii $R_{1}$, $R_{2}$ of the WarpShell metric denote inner and outer shell boundaries and are unrelated to $R_{b}$.
@@ 2.2 Hawking–Ellis classification
  • Type II: Pure radiation — stress-energy has a defective $2\times 2$ null Jordan block, a double null eigenvector with no additional linearly independent null eigenvector. Canonical example: a null dust with stress-energy $T_{ab}=\Phi^{2}k_{a}k_{b}$ for null $k^{a}$ [30].
  • Type III: $3\times 3$ null Jordan structure — a triple null eigenvector in nilpotent configuration. No known classical or semiclassical source produces Type III stress-energy; rare in physical applications [29].
@@
Our non-Type-I verification pathway (observer optimization for WEC/SEC/DEC and full null optimization for NEC) has been validated primarily on a null-dust (Type II) benchmark in the test suite; full validation on Type III and generic Jordan-block configurations remains an open direction (see Section 8).
@@
Throughout this paper, “robust margin” refers to the merged quantity (hereafter hybrid margin): algebraic slack at Type I points and $\zeta_{\max}$-capped extremum at non-Type I points. For Type I points, violation detection and the reported margins use the algebraic slack (item 2), which is cap-independent and authoritative. The observer optimizer is run additionally to provide the explicit worst-case observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$; its $\zeta_{\max}$-capped contraction minimum (item 3) is reported separately in sensitivity analyses (Table 10) but is not used for violation detection. The algebraic slack and capped extremum are in general different numerical values: the former represents the global eigenvalue-based quantity, while the latter depends on $\zeta_{\max}$. These two quantities are conceptually distinct and should not be compared directly; the algebraic slack measures distance to the violation boundary, while the capped extremum quantifies the worst-case observer-contracted energy density $T_{ab}\,u^{a}u^{b}$ within the tested observer range. The algebraic slack (e.g. $\min_{i}(\rho+p_{i})$ for NEC) certifies satisfaction or violation but is not, in general, equal to the minimum observer-contracted energy density $\min_{u}T_{ab}\,u^{a}u^{b}$; it measures proximity to the eigenvalue inequality boundary rather than a physical extremum over the observer manifold. For the metrics and rapidity caps used here ($\zeta_{\max}=5$) we empirically observe sign agreement between the algebraic Type I truth and the $\zeta_{\max}$-capped optimizer at all grid points (i.e. no violation is “hidden” above the cap). This is not guaranteed in general: a Type I tensor with $\rho+p_{i}<0$ but $|\rho+p_{i}|\ll|\rho|$ could require an arbitrarily large boost to flip the sign of the measured energy density, so a finite $\zeta_{\max}$ might in principle miss it. At non-Type I grid points (up to $\sim\!3$% for the warp metrics and <1<\!1% for WarpShell; see Table 9), the algebraic inequalities do not apply and the optimizer is the sole method; this pathway is validated via a null-dust benchmark (Section 8).
@@ 2.3 Observer parameterization
We distinguish three verification pathways. (i) For Hawking–Ellis Type I stress-energy, the eigenvalue inequalities (equations 1013) are observer-independent and provide exact certification (Section 2.2). (ii) For the NEC at non-Type I points, we optimize over the closed null cone, with directions on $S^{2}$ parameterized by $(\theta,\phi)$ via stereographic projection. (iii) For the WEC, SEC, and DEC at non-Type I points, we optimize over the open timelike cone with a rapidity cap $\zeta_{\max}$; a positive margin at the cap is a lower bound on the true margin and does not certify satisfaction for arbitrarily large boost. The phrase “over all admissible observers” refers to this capped, type-dependent pathway. We parameterize the space of timelike and null vectors at each point as follows.
@@
The rapidity parameterization has the advantage that $\zeta=0$ recovers the Eulerian observer, so the optimization always includes the Eulerian result as a baseline. The WEC and DEC involve three-dimensional optimization over $(\zeta,\theta,\phi)$, while the NEC requires only $(\theta,\phi)$. The true timelike observer manifold has unbounded rapidity ($\zeta\in[0,\infty)$); in practice, we optimize over $\zeta\in[0,\zeta_{\max}]$ with a fixed cap $\zeta_{\max}=5$ (corresponding to $\gamma\approx 74$; see Section 2.4).